Q\Determine the following ODEs are exact or not? Why?
(y+cosx)dx+(x+siny)dy=0
1
Expert's answer
2020-10-27T19:28:31-0400

"\displaystyle
(y+\cos{x})\mathrm{d}x+(x+\sin{y})\mathrm{d}y=0\\

M = y + \cos{x}\\

\frac{\partial M}{\partial y} = 1\\

N = x + \sin{y}\\

\frac{\partial N}{\partial x} = 1\\


\textsf{Therefore, since}\, \frac{\partial M}{\partial y} = \frac{\partial N}{\partial x}, \\
\textsf{the differential equation is exact.}"